# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v1

@@ 1.2. Quantum minors
Let $(a_{1},\ldots a_{m}),(b_{1},\ldots,b_{m})$ be two sequences of indices. Define a quantum minor as the sum over all permutations of $1,2,\ldots,m$, see [MolevYangians, (1.54),(1.55)]:
@@
    [MolevYangians, Proposition 1.6.8] A quantum minor can be decomposed as
@@
    [MolevYangians, Proposition 1.7.1] The following relation holds:
    $[E_{kl},T^{a_{1}\ldots a_{m}}_{b_{1}\ldots b_{m}}(u)]=\hbar\sum_{i=1}^{m}\delta_{a_{i}l}T^{a_{1}\ldots k\ldots a_{m}}_{b_{1}\ldots b_{m}}(u)-\hbar\sum_{i=1}^{m}\delta_{b_{i}k}T^{a_{1}\ldots a_{m}}_{b_{1}\ldots l\ldots b_{m}}(u),$
@@
For the last part, it is enough to consider the coefficient of $u^{0}$ on both sides of [MolevYangians, Proposition 1.7.1]. ∎
@@
By [MolevNazarovOlshanskii], the coefficients $\{A_{i}\}$ generate the center of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$.
We will need an explicit form of the inverse $T(u)^{-1}$ in terms of quantum comatrix, see [MolevYangians, Proposition 1.9.2]:
@@
where the hats indicate the omitted indices. The following equation is an immediate corollary of [MolevYangians, Definition 1.9.1] and [MolevYangians, Proposition 1.9.2]:
@@ 1.3. Differential operators
In particular, in terms of matrices from Section 1.1, we can write
@@
It is enough to show that their actions on $\mathcal{O}(\mathrm{Mat}_{N\times N})$ coincide. By the Howe duality [ChengWang, Theorem 5.16], under the $(\mathfrak{gl}_{N},\mathfrak{gl}_{N})$-action by
@@
The result follows from 1.2. ∎
@@
It follows from Proposition 7.2 and Proposition 4.7 that
@@
where the equality is understood as in Remark 4.8 and multiplication by $A^{R}(u)$ is entry-wise. Therefore, by (1.13) and $T^{R}(u)=u+E^{R}$, we have
@@
It follows from (1.14) and Proposition 1.3 that
@@ 2.1. Shibukawa-Ueno $R$-operator
In this subsection, we recall the definition of Shibukawa-Ueno $R$-operators [ShibukawaUeno] and study the induced symmetrizers. By [FelderPasquier], its elliptic version is gauge-equivalent to Belavin’s $R$-matrix that is used to define the original elliptic shuffle algebras [FeiginOdesskiiSklyanin, OdesskiiFeiginEllipticDeformation, FeiginOdesskii]. At the same time, we will show in the next subsection that its rational version is related to degenerate affine Hecke algebras.
Let $\theta(z)$ be a holomorphic function satisfying the three-term equation
(2.1) $\sum_{\sigma}\theta(u+x_{\sigma(1)})\theta(u-x_{\sigma(1)})\theta(x_{\sigma(3)}+x_{\sigma(2)})\theta(x_{\sigma(3)}-x_{\sigma(2)})=0,$
where the sum is over cyclic permutations of the set $(1,2,3)$. Its analytic solution is essentially one of the following three functions [WhittakerWatson]:
@@
where $\tau\in\mathbb{H}$ is in the upper half-plane, $\tau_{1}\in\mathbf{C}\backslash\{0\}$, and
@@
is the Jacobi’s first theta function. Observe that $\theta(-z)=-\theta(z)$ fol all of them.
Denote by
$G_{\theta}(z,w):=\frac{\theta(z+w)\theta^{\prime}(0)}{\theta(z)\theta(w)}$
@@
[ShibukawaUeno] The Shibukawa-Ueno $R$-operator associated to $\theta$ is the operator on $\mathcal{M}_{k}$ defined by
@@
[ShibukawaUeno] The Shibukawa-Ueno $R$-operator satisfies the quantum Yang-Baxter equation
@@
Consider the $\kappa$-symmetrization operator [CherednikRmatrix]
@@
Let us fix arbitrary, but pair-wise non-equal $z_{0},\ldots,z_{l}$, and consider both sides as functions of $z_{l+1}$. One can check that they are zero when $z_{l+1}=z_{i}-\kappa$ for any $1\leq i\leq l$. Indeed, since both sides are symmetric with respect to $z_{1},\ldots,z_{l}$, it is enough to show it for $z_{l+1}=z_{l}-\kappa$. The left-hand side is obviously zero, and the right-hand one becomes
$\frac{\theta(l\kappa)}{\theta(l\kappa+\kappa)}\left[\sum_{i=0}^{l-1}\prod_{\begin{subarray}{c}j=0\ldots l-1\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa)\cdot G^{\theta}(z_{l}-z_{i},l\kappa)-\prod_{i=1}^{l}G^{\theta}(z_{l}-z_{j},\kappa)\right]=0$
by induction assumption.
One can also check that both sides have simple poles at the same points with the same residues. Denote by
$\Delta(z_{l+1}):=\prod_{i=0}^{l}\theta(z_{l+1}-z_{i})\cdot\left[\prod_{i=0}^{l}G^{\theta}(z_{l+1}-z_{i},\kappa)-\sum_{i=0}^{l}\prod_{\begin{subarray}{c}j=0\ldots l\\ j\neq i\end{subarray}}G^{\theta}(z_{i}-z_{j},\kappa)\cdot G^{\theta}(z_{l+1}-z_{i},l\kappa+\kappa)\right]$
the regularized difference between the left- and the right-hand side. Then the equality of residues gives $\Delta(z_{i})=0$ for all $1\leq i\leq l$. In the elliptic case, it satisfies
$\Delta(z_{l+1}+1)=\Delta(z_{l+1}),\qquad\Delta(z_{l+1}+\tau)=(-1)^{n}e^{-(l+1)\pi i\tau+\sum_{i=0}^{l}(z_{i}-\kappa)-2\pi i(l+1)z_{l+1}}\Delta(z_{l+1}).$
In the rational case, it is a polynomial in $z_{l+1}$ of degree less or equal to $l+1$. In the trigonometric case, it is proportional to a polynomial in $\zeta_{l+1}=\exp(2\pi iz_{l+1}/\tau_{1})$ of degree less or equal to $l+1$. In all these cases, the dimension of such functions is $l+2$. Since $\Delta(z_{l+1})$ vanishes in at least $l+2$ points, it should be identically zero. ∎
@@
Let us consider $\mathrm{Sym}_{k}$ and $\mathrm{Sym}_{k}^{-}$ as elements of $\mathbf{C}[S_{k}]$ embedded into semi-direct product $\mathbf{C}[S_{k}]\ltimes\mathcal{M}_{k}$. Then Proposition 2.3 for the quantization parameter $-\kappa$ is equivalent to equality
@@ 2.2. Degenerate affine Hecke algebras and shuffle products
In this subsection, we recall some basic facts about degenerate affine Hecke algebras and relate them to Feigin-Odesskii shuffle algebras [FeiginOdesskii]. Unless otherwise stated, we refer the reader to [KirillovHecke] for all the statements and proofs (upon replacing $h$ in loc. cit. with $-\kappa$).
@@
By affinization procedure [KSQuantumGroups, Section 8.7.1], the collection of operators
@@
in $\mathbf{C}[S_{k}]$. The following is an analog of [Jucis].
@@
It is enough to check the relation in some faithful representation. As such, we can take $(\mathbf{C}^{N})^{\otimes k}$ with the usual permutation action for $N\gg 0$. Then it is the content of [MolevYangians, Proposition 1.6.2] and [MolevYangians, Proposition 1.6.3]. ∎
Consider the action of Hecke $R$-matrices in the polynomial representation (2.4). Using explicit formula (2.5) and 2.1, we see that it is essentially given by the rational Shibukawa-Ueno braiding element:
@@
The following proposition is a rational analog of the relation between Hall-Littlewood polynomials [Macdonald, Chapter III] and affine Hecke algebras (see [LusztigCharacter]), and follows immediately from Proposition 2.6, Proposition 2.3, and 2.4.
@@
with the $\mathbf{C}[\hbar]$-linear product $*^{\kappa}\colon\mathcal{F}^{\kappa}_{k}\otimes_{\mathbf{C}[\hbar]}\mathcal{F}^{\kappa}_{l}\rightarrow\mathcal{F}^{\kappa}_{k+l}$ defined by
@@
There is another deformation of the latter which is a simplest case of the Feigin-Odesskii shuffle algebras [FeiginOdesskii].
@@
with the $\mathbf{C}[\hbar]$-linear product $*^{\kappa}\colon\mathcal{S}^{\kappa}_{k}\otimes_{\mathbf{C}[\hbar]}\mathcal{S}^{\kappa}_{l}\rightarrow\mathcal{S}^{\kappa}_{k+l}$ defined by
@@
The original construction of elliptic algebras [FeiginOdesskiiSklyanin, OdesskiiFeiginEllipticDeformation, FeiginOdesskii] defines a shuffle product with factor $\vartheta_{1}(z)$ (up to shifts) on the symmetric algebra $S^{\bullet}H^{0}(\Theta_{N})$ of global sections of a degree $N$ line bundle on an elliptic curve. By [EndelmanHodgesTwistedSU], the latter admits a degeneration such that $\vartheta_{1}(z)\mapsto z$ and the symmetric algebra becomes ${}^{N}\mathcal{S}^{\kappa}$.
The following theorem is a corollary of Proposition 2.7.
@@
One can similarly relate the trigonometric Shibukawa-Ueno operator to the Laurent polynomial representation of the non-degenerate affine Hecke algebra [ChariPressley, Section 12.3] via the affinization procedure of [KSQuantumGroups]. It would be interesting to relate the elliptic Shibukawa-Ueno operators to elliptic affine Hecke algebras [GinzburgKapranovVasserot].
@@ 2.3. Shifted Yangian of $\mathfrak{sl}_{2}$
The following definition is [FrassekPestunTsymbaliuk, Section 2.1]; we use its equivalent form in terms of currents. In particular, we assume that the weight $(m_{1},m_{2})$ is antidominant, i.e. $m_{1}\leq m_{2}$.
@@
Let $m=m_{1}-m_{2}$. The definition of the shifted Yangian $\mathrm{Y}_{m}(\mathfrak{sl}_{2})$ can be found in [BravermanFinkelbergNakajima, Section B.1], see also its form in terms of currents $e(u),f(u),h(u)$ in [FKPRW, Section 6.1]. Its relation with $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ is as follows.
@@
[FrassekPestunTsymbaliuk, Proposition 2.19] There is a $\mathbf{C}[\hbar]$-algebra embedding
@@
where $f(w\pm\hbar\epsilon_{i}):=f(w_{1},\ldots,w_{i}\pm\hbar,\ldots,w_{N},\hbar)$. The homomorphism of the following theorem is called the GKLO homomorphism, see [GKLO] for the non-shifted version, [KWWY] for dominant shifts, and [BravermanFinkelbergNakajima] for general definition. We use the version adapted to $\mathfrak{gl}_{2}$.
@@
[FrassekPestunTsymbaliuk, Theorem 2.35] There is a $\mathbf{C}[\hbar]$-algebra homomorphism
@@
The following is a special case of truncated shifted Yangians, see [BravermanFinkelbergNakajima, Appendix B]
@@
Recall the shuffle algebra of 2.8. The following is [TsymbaliukPBWD, Theorem 6.20] combined with Theorem 2.11, see also [Enriquez].
@@ 3. Harish-Chandra bimodules
In what follows, we will use the following trick to relate the results of [MolevTransvectorAlgebras] to the setting of the paper. Denote by $\alpha=N+1$. Then there is an algebra map
@@ 3.1. Action of the degenerate affine Hecke algebra
By matrix form of (3.4), we need to show that $A(E)=0$. This is [MolevYangians, Theorem 7.2.1]. ∎
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\kappa}_{k}$ of 2.5. According to [Cherednik, Example 2.1], there is an action of $\mathfrak{H}^{\hbar}_{k}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$. Recall that there is an isomorphism
@@
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ from 2.5. The following can be checked directly.
@@
For the following corollary, we use the embedding of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ into $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$ as in 3.4.
@@
The action of Proposition 3.8 induces an action of $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ such that
@@
Observe that the action is well-defined since $f(\vec{Y})$ is symmetric. The spherical subalgebra $e^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}e^{\hbar}_{k}$ acts on the invariants $[\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}]^{S_{k}}$ that, by definition, can be identified with $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$. The explicit formula follows from Proposition 3.8. ∎

### v2

@@ 1.2. Quantum minors
Let $(a_{1},\ldots a_{m}),(b_{1},\ldots,b_{m})$ be two sequences of indices. Define a quantum minor as the sum over all permutations of $1,2,\ldots,m$, see [Mol07, (1.54),(1.55)]:
@@
    [Mol07, Proposition 1.6.8] A quantum minor can be decomposed as
@@
    [Mol07, Proposition 1.7.1] The following relation holds:
    $E_{kl},T^{a_{1}\ldots a_{m}}_{b_{1}\ldots b_{m}}(u)]=\hbar\sum_{i=1}^{m}\delta_{a_{i}l}T^{a_{1}\ldots k\ldots a_{m}}_{b_{1}\ldots b_{m}}(u)-\hbar\sum_{i=1}^{m}\delta_{b_{i}k}T^{a_{1}\ldots a_{m}}_{b_{1}\ldots l\ldots b_{m}}(u),$
@@
For the last part, it is enough to consider the coefficient of $u^{0}$ on both sides of [Mol07, Proposition 1.7.1]. ∎
@@
By [MNO96], the coefficients $\{A_{i}\}$ generate the center of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})$.
We will need an explicit form of the inverse $T(u)^{-1}$ in terms of quantum comatrix, see [Mol07, Proposition 1.9.2]:
@@
where the hats indicate the omitted indices. The following equation is an immediate corollary of [Mol07, Definition 1.9.1] and [Mol07, Proposition 1.9.2]:
@@ 1.3. Differential operators
In particular, in terms of matrices from Section˜1.1, we can write
@@
It is enough to show that their actions on $\mathcal{O}(\mathrm{Mat}_{N\times N})$ coincide. By the Howe duality [CW12, Theorem 5.16], under the $(\mathfrak{gl}_{N},\mathfrak{gl}_{N})$-action by
@@
The result follows from ˜1.2. ∎
@@
It follows from Proposition˜7.2 and Proposition˜4.7 that
@@
where the equality is understood as in Remark˜4.8 and multiplication by $A^{R}(u)$ is entry-wise. Therefore, by (1.13) and $T^{R}(u)=u+E^{R}$, we have
@@
It follows from (1.14) and Proposition˜1.3 that
@@ 2.1. Shibukawa-Ueno $R$-operator
In this subsection, we recall the definition of Shibukawa-Ueno $R$-operators [SU92] and study the induced symmetrizers. By [FP94], its elliptic version is gauge-equivalent to Belavin’s $R$-matrix that is used to define the original elliptic shuffle algebras [OF89, OF97, FO98]. At the same time, we will show in the next subsection that its rational version is related to degenerate affine Hecke algebras.
Let $\theta(z)$ be a skew-symmetric (i.e. $\theta(-z)=-\theta(z)$) holomorphic function satisfying the three-term relation
(2.1) $\begin{split}\theta(x+&y)\theta(x-y)\theta(z+w)\theta(z-w)+\\ +&\theta(x+z)\theta(x-z)\theta(w+y)\theta(w-y)+\\ +&\theta(x+w)\theta(x-w)\theta(y+z)\theta(y-z)=0.\end{split}$
Its analytic solution is essentially one of the following three functions [WW96]:
@@
[…]
@@
is the Jacobi’s first theta function. Denote by
$G^{\theta}(z,w):=\frac{\theta(z+w)\theta^{\prime}(0)}{\theta(z)\theta(w)}$
@@
[SU92] The Shibukawa-Ueno $R$-operator associated to $\theta$ is the operator on $\mathcal{M}_{k}$ defined by
@@
[SU92] The Shibukawa-Ueno $R$-operator satisfies the quantum Yang-Baxter equation
@@
Consider the $\kappa$-symmetrization operator [Che86]
@@
This is the content of ˜A.1 for $u=z_{l+1}$. ∎
@@
Let us consider $\mathrm{Sym}_{k}$ and $\mathrm{Sym}_{k}^{-}$ as elements of $\mathbf{C}[S_{k}]$ embedded into semi-direct product $\mathbf{C}[S_{k}]\ltimes\mathcal{M}_{k}$. Then Proposition˜2.3 for the quantization parameter $\kappa$ is equivalent to equality
@@ 2.2. Degenerate affine Hecke algebras and shuffle products
In this subsection, we recall some basic facts about degenerate affine Hecke algebras and relate them to Feigin-Odesskii shuffle algebras [FO98]. Unless otherwise stated, we refer the reader to [Kir97] for all the statements and proofs (upon replacing $h$ in loc. cit. with $\kappa$).
@@
By affinization procedure [KS97, Section 8.7.1], the collection of operators
@@
in $\mathbf{C}[S_{k}]$. The following is an analog of [Juc66].
@@
It is enough to check the relation in some faithful representation. As such, we can take $(\mathbf{C}^{N})^{\otimes k}$ with the usual permutation action for $N\gg 0$. Then it is the content of [Mol07, Proposition 1.6.2] and [Mol07, Proposition 1.6.3]. ∎
Consider the action of Hecke $R$-matrices in the polynomial representation (2.4). Using explicit formula (2.5) and ˜2.1, we see that it is essentially given by the rational Shibukawa-Ueno braiding element:
@@
The following proposition is a rational analog of the relation between Hall-Littlewood polynomials [Mac95, Chapter III] and affine Hecke algebras (see [Lus83]), and follows immediately from Proposition˜2.6, Proposition˜2.3, and ˜2.4.
@@
with the $\mathbf{C}[\hbar]$-linear product $^{\kappa}\colon\mathcal{F}^{\kappa}_{k}\otimes_{\mathbf{C}[\hbar]}\mathcal{F}^{\kappa}_{l}\rightarrow\mathcal{F}^{\kappa}_{k+l}$ defined by
@@
There is another deformation of the latter which is a simplest case of the Feigin-Odesskii shuffle algebras [FO98].
@@
with the $\mathbf{C}[\hbar]$-linear product $^{\kappa}\colon\mathcal{S}^{\kappa}_{k}\otimes_{\mathbf{C}[\hbar]}\mathcal{S}^{\kappa}_{l}\rightarrow\mathcal{S}^{\kappa}_{k+l}$ defined by
@@
The original construction of elliptic algebras [OF89, OF97, FO98] defines a shuffle product with factor $\vartheta_{1}(z)$ (up to shifts) on the symmetric algebra $S^{\bullet}H^{0}(\Theta_{N})$ of global sections of a degree $N$ line bundle on an elliptic curve. By [EH04], the latter admits a degeneration such that $\vartheta_{1}(z)\mapsto z$ and the symmetric algebra becomes $^{N}\mathcal{S}^{\kappa}$.
The following theorem is a corollary of Proposition˜2.7.
@@
One can similarly relate the trigonometric Shibukawa-Ueno operator to the Laurent polynomial representation of the non-degenerate affine Hecke algebra [CP95, Section 12.3] via the affinization procedure of [KS97]. It would be interesting to relate the elliptic Shibukawa-Ueno operators to elliptic affine Hecke algebras [GKV97].
@@ 2.3. Shifted Yangian of $\mathfrak{sl}_{2}$
The following definition is [FPT22, Section 2.1]; we use its equivalent form in terms of currents. In particular, we assume that the weight $(m_{1},m_{2})$ is antidominant, i.e. $m_{1}\leq m_{2}$.
@@
Let $m=m_{1}-m_{2}$. The definition of the shifted Yangian $\mathrm{Y}_{m}(\mathfrak{sl}_{2})$ can be found in [BFN19, Section B.1], see also its form in terms of currents $e(u),f(u),h(u)$ in [Fin+18, Section 6.1]. Its relation with $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ is as follows.
@@
[FPT22, Proposition 2.19] There is a $\mathbf{C}[\hbar]$-algebra embedding
@@
where $f(w\pm\hbar\epsilon_{i}):=f(w_{1},\ldots,w_{i}\pm\hbar,\ldots,w_{N},\hbar)$. The homomorphism of the following theorem is called the GKLO homomorphism, see [Ger+05] for the non-shifted version, [Kam+14] for dominant shifts, and [BFN19] for general definition. We use the version adapted to $\mathfrak{gl}_{2}$.
@@
[FPT22, Theorem 2.35] There is a $\mathbf{C}[\hbar]$-algebra homomorphism
@@
The following is a special case of truncated shifted Yangians, see [BFN19, Appendix B]
@@
Recall the shuffle algebra of ˜2.8. The following is [Tsy21, Theorem 6.20] combined with Theorem˜2.11, see also [Enr03].
@@ 3. Harish-Chandra bimodules
In what follows, we will use the following trick to relate the results of [Mol02] to the setting of the paper. Denote by $\alpha=N+1$. Then there is an algebra map
@@ 3.1. Action of the degenerate affine Hecke algebra
By matrix form of (3.4), we need to show that $A(E)=0$. This is [Mol07, Theorem 7.2.1]. ∎
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\kappa}_{k}$ of ˜2.5. According to [Che98, Example 2.1], there is an action of $\mathfrak{H}^{\hbar}_{k}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}$. Recall that there is an isomorphism
@@
Recall the degenerate affine Hecke algebra $\mathfrak{H}^{\hbar}_{k}$ from ˜2.5. The following can be checked directly.
@@
For the following corollary, we use the embedding of $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ into $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\ltimes S^{\bullet}\mathbf{C}^{N}$ as in ˜3.4.
@@
The action of Proposition˜3.8 induces an action of $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}$ on the Harish-Chandra bimodule $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$ such that
@@
Observe that the action is well-defined since $f(\vec{Y})$ is symmetric. The spherical subalgebra $e^{\hbar}_{k}\mathfrak{H}^{\hbar}_{k}e^{\hbar}_{k}$ acts on the invariants $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes(\mathbf{C}^{N})^{\otimes k}]^{S_{k}}$ that, by definition, can be identified with $\mathrm{U}_{\hbar}(\mathfrak{gl}_{N})\otimes S^{k}\mathbf{C}^{N}$. The explicit formula follows from Proposition˜3.8. ∎